# Large Language Models in Bioinformatics: A Survey

### v1

@@ 1 Introduction
Bioinformatics is an interdisciplinary field that combines biology, computer science, and information technology to analyze and interpret complex biological data Abdi et al. (2024). Recently, LLMs have demonstrated remarkable progress in the domain of natural language processing (NLP), with applications that span a wide array of tasks Min et al. (2023); Raiaan et al. (2024). However, the nature of biological data and the associated tasks differ significantly from text data, presenting unique challenges. The accurate and precise handling of biomedical data to effectively form features and embeddings suitable for LLMs is an ongoing challenge that necessitates innovative solutions.
Within the biological domain, tasks exhibit a high degree of variability and specificity. These include the functional prediction and generation of DNA sequences, the prediction of RNA structure and function, the prediction and design of protein structures, and the analysis of single-cell data, which encompasses dimensionality reduction, clustering, cell annotation, and developmental trajectory analysis. There is a growing interest among researchers in harnessing the power of LLMs for bioinformatics and computational biology, yielding significant results. As illustrated in Figure1, the development, training, and application of large models in bioinformatics are increasing at a rapid pace. Despite this, the diverse methodologies focused on these various tasks have not been systematically summarized and analyzed, presenting an opportunity for comprehensive review and synthesis.
Organization of This Survey: This paper reviews recent advancements in LLMs for bioinformatics. We begin with preliminary concepts (§2), covering key architectures and their relevance to biological data. Next, an overview of representative LLMs in bioinformatics is presented in Table1. We then explore LLM-driven innovations across DNA and Genomics (§3), RNA (§4), Proteins (§5), and Single-Cell Analysis (§6). Finally, we discuss key challenges (§7.1) and propose future directions (§7.2), emphasizing multimodal learning, hybrid AI models, and clinical applications. To conclude, we analyze the limitations of this survey, highlighting areas that require further exploration to fully capture the evolving landscape of LLMs in bioinformatics.
Figure 1: Overview of LLM advancements and applications in bioinformatics, spanning DNA, RNA, protein, and scRNA.
@@ 2 Preliminaries
Encoder-decoder models, such as T5-based and transformer-based architectures (e.g., RoseTTAFold Baek et al. (2021)), are designed for sequence-to-sequence tasks, where an input sequence is transformed into an output sequence. This architecture is particularly useful for the tasks involving mapping between different biological modalities, (e.g., gene expression predication, multiomics data integration). For RoseTTAFold, it employs a three-track neural network to predict protein interactions as well as complex formations. Meanwhile, the encoder-decoder architecture shows great potential when applied to the tasks that require bidirectional context understanding and structured output generation, represented by RNA secondary structure prediction (e.g., RhoFold+ Shen et al. (2024)) and genome-wide variant effect prediction. However, these models often require substantial computational resources for both training and inference, making them less accessible for researchers with limited computational infrastructure. Additionally, their performance is highly dependent on large-scale domain-specific pre-training, necessitating extensive datasets for generalization.
@@
AlphaFold2 Jumper et al., 2021 Nature - PDB, BFD, etc Protein Predicts protein 3D structures from sequences
@@
RoseTTAFold Minkyung Baek et al., 2021 Science Encoder-decoder PDB, etc Protein Predicts protein structures and interactions efficiently
@@
Enformer Avsec, Ž et al., 2021 Nat Methods Encoder-only UCSC Genome Browser, etc DNA Predicts gene expression from DNA sequences
@@
ESM-1b A. Rives, J et al., 2021 PNAS Encoder-only UniParc Protein Predicts protein structure and function
@@
MSA Transformer Roshan M Rao et al., 2021 PMLR Encoder-only UniRef50 Protein Predicts protein structures using multiple sequence alignments
@@
Fold2Seq Yue Cao et al., 2021 PMLR Encoder-decoder CATH 4.2 Protein Designs protein sequences from 3D folds
@@
ProGPT2 Ferruz, N et al., 2022 Nat Commun Decoder-only Uniref50 Protein Generates novel protein sequences with natural-like properties
@@
scBERT Yang, F et al., 2022 Nat Mach Intell Encoder-only Panglao scRNA Annotates cell types from single-cell RNA-seq data
@@
ProteinBERT Nadav Brandes et al., 2022 Bioinformatics Encoder-only UniProtKB and UniRef90 Protein Predicts protein functions from sequences
@@
ProtTrans A. Elnaggar et al., 2022 IEEE Encoder-only UniRef and BFD Protein Predicts protein functions and structures
@@
ProGen Madani, A et al., 2023 Nat Biotechnol Decoder-only UniprotKB, etc Protein Generates functional protein sequences across families
@@
ESMFold Zeming Lin et al., 2023 Science Encoder-decoder UniRef50 Protein Predicts protein structures from single sequences
@@
Geneformer Theodoris, C.V. et al., 2023 Nature Encoder-decoder GEO, SRA, HCA, etc scRNA Predicts gene networks from single-cell data
@@
HyenaDNA Eric Nguyen et al., 2023 NeurIPS Decoder-only Human reference genome DNA Models long-range DNA interactions at single-nucleotide resolution
@@
Uni-RNA Xi Wang et al., 2023 bioRxiv Encoder-only RNAcentral, etc RNA Predicts RNA structures, functions, and properties
@@
ProGen2 Erik Nijkamp et al., 2023 Cell Systems Decoder-only UniProtKB and BFD Protein Generates functional protein sequences across families
@@
xTrimoPGLM Bo Chen et al., 2023 arXiv Encoder-decoder Uniref90 and ColabFoldDB Protein Predicts and designs protein sequences and structures
@@
RNA-MSM Yikun Zhang et al., 2023 Nucleic Acids Research Encoder-only RNAcmap RNA Predicts RNA structures using evolutionary information
@@
TFBert Luo H et al., 2023 Interdiscip Sci Encoder-only 690 ChIP-seq DNA Predicts transcription factor binding sites
@@
DNAGPT Daoan Zhang et al., 2023 arXiv Decoder-only GRCh38 DNA Generates and analyzes DNA sequences
@@
GROVER Sanabria et al., 2024 Nat Mach Intell Encoder-only GRCh37 DNA Predicts DNA functions from sequence context
@@
megaDNA Shao, B et al., 2024 Nat Commun Decoder-only NCBI GenBank DNA Generates and analyzes functional genomes
@@
Nucleotide Transformer Dalla-Torre, H et al., 2024 Nat Methods Encoder-only GRCh38 DNA Predicts molecular phenotypes from DNA sequences
@@
RNAErnie Wang, N et al., 2024 Nat Mach Intell Encoder-only RNAcentral RNA Predicts RNA functions and structures
@@
RhoFold+ Shen, T et al., 2024 Nat Methods Encoder-only RNAcentral RNA Predicts RNA 3D structures from sequences
@@
GPTCelltype Hou, W et al., 2024 Nat Methods Decoder-only figshare, Zenodo, GEO, etc. scRNA Automates cell type annotation using GPT-4
@@
Evo Eric Nguyen et al., 2024 Science Decoder-only OpenGenome DNA Predicts and designs DNA, RNA, proteins
@@
scFoundation Hao, M et al., 2024 Nat Methods Encoder-decoder HCA, Single Cell Portal, GEO, etc scRNA Predicts and analyzes single-cell transcriptomics data
@@
scGPT Cui, H et al., 2024 Nat Methods Encoder-decoder CELLxGENE scRNA Predicts and analyzes single-cell omics data
@@
AlphaFold3 Abramson, J., 2024 Nature - PDB Protein Predicts biomolecular structures and interactions accurately
@@
DNABERT2 Zhihan Zhou et al., 2024 ICLR Encoder-only Human and multi-species genome DNA Predicts genomic functions across species efficiently
@@
ESM-DBP Zeng, W et al., 2024 Nat Commun Encoder-only UniProtKB Protein Predicts DNA-binding proteins and residues accurately
@@
RoseTTAFold All-Atom Rohith Krishna et al., 2024 Science Encoder-decoder PDB, etc Protein Predicts and designs biomolecular structures
@@
ProstT5 Michael Heinzinger et al., 2024 NAR Genomics and Bioinformatics Encoder-decoder AFDB Protein Translates protein sequences to 3D structures
@@
EpiGePT Gao, Z et al., 2024 Genome Biol Encoder-only ENCODE DNA Predicts context-specific epigenomic signals and interactions
@@
RiNALMo Rafael Josip Penić et al., 2024 arXiv Encoder-only RNAcentral RNA Predicts RNA structures and functions
@@
ENBED Aditya Malusare et al., 2024 Bioinformatics Advances Encoder-decoder NCBI-Genome DNA Analyzes DNA sequences with byte-level precision
@@
GPN-MSA Benegas, G et al., 2025 Nat Biotechnol Encoder-only multiz MSA, etc DNA Predicts genome-wide variant effects efficiently
@@
ESM-3 Thomas Hayes et al., 2025 Science Encoder-decoder UniRef Protein Predicts and designs proteins with multi-modal inputs
Table 1: Overview of representative LLMs in bioinformatics, categorized by architecture, dataset, task, and application domain.
@@ 3 DNA and Genomics: Learn and Generate
DNABERT-2 is a high-performance foundation model based on the Transformer architecture, designed for multi-species genome analysis. By incorporating innovative tokenization methods, efficient attention mechanisms, and a comprehensive evaluation benchmark, it provides a powerful tool for understanding and interpreting genomic data across diverse species Feng et al. (2024).
@@ 4.1.1 RNA secondary structure prediction
Some researchers benchmarked 6 RNA-LLMs which include RNABERT, RNA-FM, RNA-MSM, ERNIE-RNA, RNAErnie and RiNALMo for their ability to predict RNA secondary structure. They found that RiNALMo and ERNIE-RNA were the models that could better represent and separate the RNA families in the projection without almost overlap. Zablocki et al. (2024)
@@ 4.1.2 RNA Teridry structure prediction
Uni-RNA represents a paradigm shift in RNA research by combining large-scale pre-training with advanced deep learning techniques. Its ability to accurately predict RNA structures, functions, and properties positions it as a powerful tool for accelerating discoveries in RNA biology and therapeutic development Wang et al. (2023).
@@ 4.2 RNA Function
BEACON comprises 13 distinct tasks derived from extensive prior research, covering structural analysis, functional studies, and engineering applications.Functional tasks focus on the biological roles of RNA, including splice site prediction and non-coding RNA function classification.
RNA-RNA interactions are involved in post-transcriptional processes, contributing to gene expression regulation. RNA-protein interactions are vital for maintaining cellular homeostasis, and disruptions in these interactions can lead to cellular dysfunctions or diseases such as cancer. Furthermore, RNA-small molecule interactions have significant implications in therapeutic development, as RNAs can serve as potential drug targets, especially when conventional protein targets are less accessible.
@@ 5 Protein: Prediction and Design
Recently, LLMs have emerged as transformative computational tools in protein research, demonstrating remarkable potential to advance both fundamental comprehension and applied engineering of protein structures and functions. In this section, we classify protein-related LLMs into encoder-only, decoder-only and encoder-decoder architectures based on their diverse protein research applications, such as protein structure and function prediction and protein generation and design.
@@ 5.1 Protein Structure and Function
ESM-1b leverages a transformer-based encoder architecture to infer protein tertiary structures and functional characteristics through self-supervised learning on large-scale protein sequence databases, without relying on manual annotations of the sequenceMeier et al. (2021).

### v2

@@ 1 Introduction
Bioinformatics is an interdisciplinary field that combines biology, computer science, and information technology to analyze and interpret complex biological data Lu et al. (2020); Abdi et al. (2024); Li et al. (2024); Jiang et al. (2025b). Recently, Large Language Models (LLMs) have demonstrated remarkable progress in the domain of natural language processing (NLP), with applications that span a wide array of tasks Min et al. (2023); Jiang et al. (2023); Raiaan et al. (2024); Jiang et al. (2025a). However, the nature of biological data and the associated tasks differ significantly from text data, presenting unique challenges. The accurate and precise handling of biomedical data to effectively form features and embeddings suitable for LLMs is an ongoing challenge that necessitates innovative solutions Chen et al. (2022); Wang et al. (2023a); ESM Team (2024); Liu et al. (2025).
Within the biological domain, tasks exhibit a high degree of variability and specificity. These include the functional prediction and generation of DNA sequences Nguyen et al. (2024), the prediction of RNA structure and function Chen et al. (2022); Shen et al. (2024), the prediction and design of protein structures Jumper et al. (2020, 2021); Team (2024), and the analysis of single-cell data Wolf et al. (2018), which encompasses dimensionality reduction, clustering, cell annotation, and developmental trajectory analysis. There is a growing interest among researchers in harnessing the power of LLMs for bioinformatics and computational biology, yielding significant results Ferruz et al. (2022); Nijkamp et al. (2023); Li et al. (2025). As illustrated in Figure 1, the development, training, and application of large models in bioinformatics are increasing at a rapid pace. Despite this, the diverse methodologies focused on these various tasks have not been systematically summarized and analyzed, presenting an opportunity for comprehensive review and synthesis.
Figure 1: Selected milestones in the integration of LLMs into bioinformatics: Advances in applications to DNA, RNA, protein, and single-cell RNA (scRNA).
Organization of This Survey: This paper provides a comprehensive review of recent advancements in LLMs for bioinformatics. We begin with preliminary concepts (§2), covering key model architectures and their relevance to biological data. Next, an overview of representative LLMs applied in bioinformatics is presented in Table 1, followed by Table 3, which showcases a quantitative profile of computational costs associated with training these LLMs. Table 2 further synthesizes statistics derived from the surveyed models, highlighting average training duration and graphical memory usage across different model types. Subsequent sections explore LLM-driven innovations across various subdomains, DNA and Genomics (§3), RNA (§4), Proteins (§5), and Single-Cell Analysis (§6). Finally, we discuss key challenges (§7.1) and propose future directions (§7.2), emphasizing multimodal learning, hybrid AI models, and clinical applications. To conclude, we analyze the limitations of this survey, highlighting areas that require further exploration to fully capture the evolving landscape of LLMs in bioinformatics.
@@ 2.3 Encoder-Decoder
Encoder-decoder models, such as T5-based and transformer-based architectures (e.g., RoseTTAFold Baek et al. (2021)), are designed for sequence-to-sequence tasks, where an input sequence is transformed into an output sequence. This architecture is particularly useful for the tasks involving mapping between different biological modalities, (e.g., gene expression predication, multiomics data integration). For RoseTTAFold, it employs a three-track neural network to predict protein interactions as well as complex formations. Meanwhile, the encoder-decoder architecture shows great potential when applied to the tasks that require bidirectional context understanding and structured output generation, represented by RNA secondary structure prediction (e.g., RhoFold+ Shen et al. (2024)) and genome-wide variant effect prediction. However, these models often require substantial computational resources for both training and inference, making them less accessible for researchers with limited computational infrastructure (Table 3 and Table 2). Additionally, their performance is highly dependent on large-scale domain-specific pre-training, necessitating extensive datasets for generalization.
@@
AlphaFold2 Jumper et al. (2021) Nature PDB, BFD, etc Protein Predicts protein 3D structures from sequences
@@
RoseTTAFold Baek et al. (2021) Science Encoder-Decoder PDB, etc Protein Predicts protein structures and interactions efficiently
@@
Enformer Avsec et al. (2021) Nat. Methods Encoder-only UCSC Genome Browser, etc DNA Predicts gene expression from DNA sequences
@@
ESM-1b Rives et al. (2021) PNAS Encoder-only UniParc Protein Predicts protein structure and function
@@
MSA Transformer Rao et al. (2021) PMLR Encoder-only UniRef50 Protein Predicts protein structures using multiple sequence alignments
@@
Fold2Seq Cao et al. (2021) PMLR Encoder-Decoder CATH 4.2 Protein Designs protein sequences from 3D folds
@@
ProGPT2 Ferruz et al. (2022) Nat. Commun. Decoder-only Uniref50 Protein Generates novel protein sequences with natural-like properties
@@
scBERT Yang et al. (2022) Nat. Mach. Intell. Encoder-only Panglao scRNA Annotates cell types from single-cell RNA-seq data
@@
ProteinBERT Brandes et al. (2022) Bioinformatics Encoder-only UniProtKB and UniRef90 Protein Predicts protein functions from sequences
@@
ProtTrans Elnaggar et al. (2021) IEEE PAMI Encoder-only UniRef and BFD Protein Predicts protein functions and structures
@@
ProGen Madani et al. (2023) Nat. Biotechnol. Decoder-only UniprotKB, etc Protein Generates functional protein sequences across families
@@
ESMFold Lin et al. (2023) Science Encoder-Decoder UniRef50 Protein Predicts protein structures from single sequences
@@
Geneformer Theodoris et al. (2023) Nature Encoder-Decoder GEO, SRA, HCA, etc scRNA Predicts gene networks from single-cell data
@@
HyenaDNA Nguyen et al. (2023) NeurIPS Decoder-only Human reference genome DNA Models long-range DNA interactions at single-nucleotide resolution
@@
Uni-RNA Wang et al. (2023b) bioRxiv Encoder-only RNAcentral, etc RNA Predicts RNA structures, functions, and properties
@@
ProGen2 Nijkamp et al. (2023) Cell Syst. Decoder-only UniProtKB and BFD Protein Generates functional protein sequences across families
@@
xTrimoPGLM Chen et al. (2025) Nat. Methods Encoder-Decoder Uniref90 and ColabFoldDB Protein Predicts and designs protein sequences and structures
@@
RNA-MSM Zhang et al. (2024) Nucleic Acids Res. Encoder-only RNAcmap RNA Predicts RNA structures using evolutionary information
@@
TFBert Luo et al. (2023) Interdiscip. Sci. Comput. Life Sci. Encoder-only 690 ChIP-seq DNA Predicts transcription factor binding sites
@@
DNAGPT Zhang et al. (2023) arXiv Decoder-only GRCh38 DNA Generates and analyzes DNA sequences
@@
GROVER Sanabria et al. (2024) Nat. Mach. Intell. Encoder-only GRCh37 DNA Predicts DNA functions from sequence context
@@
megaDNA Shao and Yan (2024) Nat. Commun. Decoder-only NCBI GenBank DNA Generates and analyzes functional genomes
@@
Nucleotide Transformer Dalla-Torre et al. (2024) Nat. Methods Encoder-only GRCh38 DNA Predicts molecular phenotypes from DNA sequences
@@
RNAErnie Wang et al. (2024) Nat. Mach. Intell. Encoder-only RNAcentral RNA Predicts RNA functions and structures
@@
RhoFold+ Shen et al. (2024) Nat. Methods Encoder-only RNAcentral RNA Predicts RNA 3D structures from sequences
@@
GPTCelltype Hou and Ji (2024) Nat. Methods Decoder-only figshare, Zenodo, GEO, etc. scRNA Automates cell type annotation using GPT-4
@@
Evo Nguyen et al. (2024) Science Decoder-only OpenGenome DNA Predicts and designs DNA, RNA, proteins
@@
scFoundation Hao et al. (2024a) Nat. Methods Encoder-Decoder HCA, Single Cell Portal, GEO, etc scRNA Predicts and analyzes single-cell transcriptomics data
@@
scGPT Cui et al. (2024) Nat. Methods Encoder-Decoder CELLxGENE scRNA Predicts and analyzes single-cell omics data
@@
AlphaFold3 Abramson et al. (2024) Nature PDB Protein Predicts biomolecular structures and interactions accurately
@@
DNABERT2 Zhou et al. (2023) ICLR Encoder-only Human and multi-species genome DNA Predicts genomic functions across species efficiently
@@
ESM-DBP Zeng et al. (2024) Nat. Commun. Encoder-only UniProtKB Protein Predicts DNA-binding proteins and residues accurately
@@
RoseTTAFold All-Atom Krishna et al. (2024) Science Encoder-Decoder PDB, etc Protein Predicts and designs biomolecular structures
@@
ProstT5 Heinzinger et al. (2024) NAR Genom. Bioinform. Encoder-Decoder AFDB Protein Translates protein sequences to 3D structures
@@
EpiGePT Gao et al. (2024) Genome Biol. Encoder-only ENCODE DNA Predicts context-specific epigenomic signals and interactions
@@
RiNALMo Penić et al. (2024) arXiv Encoder-only RNAcentral RNA Predicts RNA structures and functions
@@
ENBED Malusare et al. (2024) Bioinform. adv. Encoder-Decoder NCBI-Genome DNA Analyzes DNA sequences with byte-level precision
@@
GPN-MSA Benegas et al. (2025) Nat. Biotechnol. Encoder-only multiz MSA, etc DNA Predicts genome-wide variant effects efficiently
@@
ESM-3 Hayes et al. (2025) Science Encoder-Decoder UniRef Protein Predicts and designs proteins with multi-modal inputs
Table 1: Comprehensive overview of representative LLMs in bioinformatics, systematically categorized by model architecture, dataset, task, and application domain. This table summarizes key advances in LLM-driven methods across major bioinformatics modalities, including DNA, RNA, protein, and single-cell omics. Dashes ("") indicate information not applicable for the respective model.
@@ 3 DNA and Genomics: Learn and Generate
DNABERT-2 is a high-performance foundation model based on the Transformer architecture, designed for multi-species genome analysis Zhou et al. (2023). By incorporating innovative tokenization methods, efficient attention mechanisms, and a comprehensive evaluation benchmark, it provides a powerful tool for understanding and interpreting genomic data across diverse species Feng et al. (2024).
@@ 4.1.1 RNA secondary structure prediction
Some researchers benchmarked 6 RNA-LLMs which include RNABERT, RNA-FM, RNA-MSM, ERNIE-RNA, RNAErnie and RiNALMo for their ability to predict RNA secondary structure. They found that RiNALMo and ERNIE-RNA were the models that could better represent and separate the RNA families in the projection without almost overlap Zablocki et al. (2024).
@@ 4.1.2 RNA Teridry structure prediction
Uni-RNA represents a paradigm shift in RNA research by combining large-scale pre-training with advanced deep learning techniques. Its ability to accurately predict RNA structures, functions, and properties positions it as a powerful tool for accelerating discoveries in RNA biology and therapeutic development Wang et al. (2023b).
@@ 4.2 RNA Function
BEACON comprises 13 distinct tasks derived from extensive prior research, covering structural analysis, functional studies, and engineering applications Ren et al. (2024). Functional tasks focus on the biological roles of RNA, including splice site prediction and non-coding RNA function classification.
RNA-RNA interactions are involved in post-transcriptional processes, contributing to gene expression regulation Wang et al. (2022). RNA-protein interactions are vital for maintaining cellular homeostasis Gallardo-Dodd and Kutter (2024), and disruptions in these interactions can lead to cellular dysfunctions or diseases such as cancer Liu et al. (2020). Furthermore, RNA-small molecule interactions have significant implications in therapeutic development, as RNAs can serve as potential drug targets, especially when conventional protein targets are less accessible Kaur et al. (2024).
@@ 5 Protein: Prediction and Design
Recently, LLMs have emerged as transformative computational tools in protein research, demonstrating remarkable potential to advance both fundamental comprehension and applied engineering of protein structures and functions. In this section, we introduce representative protein-related LLMs, as categorized in Table 1, based on their model architectures and diverse protein research applications, such as protein structure and function prediction and protein generation and design.
@@ 5.1 Protein Structure and Function
ESM-1b leverages a transformer-based encoder architecture to infer protein tertiary structures and functional characteristics through self-supervised learning on large-scale protein sequence databases, without relying on manual annotations of the sequence Rives et al. (2021); Meier et al. (2021).